Filtering again – University of Copenhagen

11 July 2018

 Wednesday 11th of July 2018

Filtering operations at EGRIP today. From left to right: Barbara and Jan T presents slush recovered from the hole, while Jan J stands beside them after working on winch controller settings. Night shift Vasileios and Julian in the operators’ cabin and Steff and Nan inspecting another batch of slush.

Filtering again

The saga of the chips unfortunately seems hard to complete. Despite yesterday’s good performance of the deep drill here at EGRIP, the bore hole was again found to be full of fine chips, which are hard to recover. Hence all day went with filtering the bore hole today. Julian and Vasileios now forms a filtering night shift in an attempt to keep the hole clean by removing chips every day. Removing two persons from the processing line will of course slow operations there, but for now the processing does not seem to be the bottle neck in our operations here at EGRIP. At RADIX the fight with the chips is also ongoing, here a mix of casing material and ice seems to prevent drill penetration below the casing.
The crews in the processing line the isotope CFA lab. and the Physical Properties labs. all continued making good progress during the day, cutting and measuring samples at a steady pace. The processing team has now started cutting and measuring more of this year’s deep brittle ice, as it can now be done without compromising core quality. Preparations for the next flight period continues with grooming of the skiway during the entire night.

What we did today:

  1. Drilling of 6.9 m of ice core. Filtering: 98 kg of slush recovered.
  2. Logging of main core, depth at 19.00: 1398.27 m.
  3. Processing in the science trench, final depth: 1373.35 m. Brittle ice 1062.05 m to 1081.3 m.
  4. Measurements in isotope laboratory, core drilled in 2017 completed (at 904.75 m). 2018 core: 1140.70 m to 1147.30 m.
  5. Measurements in physical properties laboratory of volume samples and brittle ice sections.
  6. Water vapour sampling and measuring.
  7. Japanese snow crystal and surface monitoring.
  8. Drilling more tiny pieces of ice and casing at the RADIX site.
  9. Test of Doppler radar, expansion of alarm-free zone out to 900m from Dome.
  10. Grooming of skiway.

Weather: Sunny most of the day, with few high clouds. Temperature -20 °C to -10 °C. Wind mainly from WSW up to 15kt.

FL, Bo Vinther

← Previous entry   Next entry →